SYSTEMATICS IN THE IR ABSORPTIONS OF ODD CARBON CHAINS

\underline{DMITRY STRELNIKOV}, ROMAN REUSCH, and WOLFGANG KRÄTSCHMER, Max-Planck-Institut für Kernphysik, D-69029 Heidelberg, PO Box 103980, Germany.

The selective laser induced oxidation method was applied to identify the IR absorptions of long carbon chain molecules isolated in oxygen containing cryogenic matrices. We have found a regular pattern in the absorptions of C_{2n+1} (n≥4) in which with increasing chain length the most intense IR absorption shifts to the red. In addition, our data allow conclusions regarding the spectra of unknown carbon chain oxides.